OpenStudy (47shift):

Ponce Leon has a principal of $900 in his savings account on October 1. The money earns an APR of 6.5% calculated quarterly as simple interest. What is the amount in the account on July 1 of the following year?

OpenStudy (47shift):

any help..?

OpenStudy (47shift):

what great help you guys have been. i adore your passion in helping me find the answer to this question. smfh

OpenStudy (zpupster):

A = P(1 + rt) is a simple interest formula A=amount P = principal = 900 r = rate = 6.5 t = time= 3/4 or 9 months/ 12 months plugging in A = 900(1 + (0.06 × 0.75))